OpenStudy (anonymous):

where do i insert parentheses to make 5.8 = 6*6-7/5*5

OpenStudy (debbieg):

Are you sure that you have copied the problem down exactly correctly? I'm having trouble getting it to work.

OpenStudy (debbieg):

5.8=5+4/5=29/5 in lowest form. Looking at what you have to work with, there is only one place involving division, and the choices there are {something}/5 if left as-is or put a ")" after that first 5, or {something}/25, if you put ( ) around the 5*5 at the end. In the latter case, that means you need to get the rest of what is give to =(29*5) for the numerator. All you have to work with is 6*6-7, which is 29, but then you get (29/25) which is not=6.8 by a factor of 5. It seems like you are missing a factor of 5??

OpenStudy (anonymous):

Here is the picture of the question from the math text book. Question #12 part b

OpenStudy (debbieg):

Wow, ok.... so what it says is \[\Large \frac{ 6^2-7 }{5 }\neq6\cdot6-7/5\cdot5\] which is certainly true. Without that weird final "5" on the RHS, you would have: \[\Large \frac{ 6^2-7 }{5 }\neq6\cdot6-7/5=6^2-\frac{ 7 }{ 5 }\]and the problem would make a whole lot more sense. Then you could simply put in the parens like so: \[\Large \frac{ 6^2-7 }{5 }=(6\cdot6-7)/5\]to make them equal. In my opinion that last *5 is a typo...??

OpenStudy (anonymous):

Thanks!

OpenStudy (debbieg):

You're welcome. :)